OpenStudy (anonymous):

i^(3/4) in rectangular form????

myininaya (myininaya):

first write i in polar form

OpenStudy (anonymous):

how do i do that?): @myininaya

myininaya (myininaya):

so you want to write i in this form: \[\cos(\theta)+i \sin(\theta)\] basically you want to find theta such that expression is equal to i

myininaya (myininaya):

can you graph i on the rectangular coordinate system?

myininaya (myininaya):

and tell me what theta is there

OpenStudy (anonymous):

After expressing it in rectangular form, you use De Moivre's theorem :)